### v1

# http://www.michaeltimothybennett.com/

Abstract
Simplicity is held by many to be the key to general intelligence. Simpler models tend to “generalise”, identifying the cause or generator of data with greater sample efficiency. The implications of the correlation between simplicity and generalisation extend far beyond computer science, addressing questions of physics and even biology. Yet simplicity is a property of form, while generalisation is of function. In interactive settings, any correlation between the two depends on interpretation. In theory there could be no correlation and yet in practice, there is. Previous theoretical work showed generalisation to be a consequence of “weak” constraints implied by function, not form. Experiments demonstrated choosing weak constraints over simple forms yielded a $110-500\%$ improvement in generalisation rate. Here we show that if one does not presuppose an abstraction layer, then all forms have equal complexity. However, in the context of a spatially and temporally extended abstraction layer, efficiency demands weak constraints take simple forms, and simplicity becomes correlated with generalisation. Simplicity has no causal influence on generalisation, but appears to due to confounding.
1 Introduction
Complexity is a quality of systems we find difficult to understand. Formal analogues include entropy [1], compression [2, 3] and even fractal dimension [4]. Physicist Leonard Susskind believes complexity may be the key to a unified theory of physics [5, 6]. Cyberneticist Francis Heylighen recently argued that goals are attractors of dynamical systems that self organise in complex reaction networks [7]. If complex reaction networks do self organise as he argues [8], then it goes some way towards explaining the origins of goal directed behaviour, and thus life. Finally, many hold that complexity is the key to general intelligence. Language models like GPT-4 amount to compressed representations of human language [9]. Simpler objects can be compressed to greater extents, because they exhibit self similarity. As a result, some hold that compression is general intelligence, meaning a general reinforcement learning agent like AIXI [10] can use Solomonoff Induction [11] to maximise expected reward across a wide range of environments. Yet for all that complexity seems to be at the heart of every matter, it has profound flaws. As a qualitative indicator of how subjectively difficult a system might be to understand, it makes perfect sense. It makes far less as an indicator of anything objective. Ockham’s Razor, the epistemic principle that simpler statements are more likely to hold true, can be understood as the claim that our subjective perceptions of complexity reflect an objective property of our environment. There is no obvious reason this should be the case, and yet it is [12]. Simpler statements tend to be more accurate representations of reality than complex ones. The aforementioned Solomonoff Induction formalises Ockham’s Razor, meaning AIXI is based on the premise that simpler models are more accurate representations of the environment. AIXI is a superintelligence in the sense that it maximises Legg-Hutter intelligence [13]. However, Jan Leike later showed that any claim regarding AIXI’s performance is “entirely subjective”. Legg-Hutter intelligence is measured with respect to a fixed Universal Turning Machine (UTM), and AIXI is only optimal if it uses exactly the same UTM [14]. This calls into question the viability of complexity based induction systems in interactive settings. Their performance is subjective, and from a pragmatic standpoint it is only objective performance claims that matter. Leike’s result suggests there could be no correlation between objective performance and subjective complexity. This concurs with what seems intuitively obvious, that complexity is an aspect of interpretation. Complexity is a measure of form, not function. So why does the subjective perception of simplicity tend to correlate with objective performance?
1.0.1 What exactly is complexity supposed to indicate?:
As it is used in AGI research [10], complexity is intended to help us infer the program which generated or caused data. If one can identify that which caused past data, then one can “generalise” to predict the outcomes in future interactions, to maximise performance [15]. We are concerned with adaptation or “the ability to generalise” [16], not any specific circumstance. This is because any system, even a rote learner, can eventually identify cause given enough data and memory (by simply rote learning every outcome until it has a complete behavioural specification of the causal program). So assuming one can correctly infer cause, then the amount of data one requires to do so is the sole measure of performance. We refer to this as sample efficiency. The more sample efficiently one can infer cause, the greater one’s ability to generalise and adapt to any desired end. Thus, we take intelligence to be a measure of the sample efficiency in generalisation.
1.0.2 Key questions:
We build upon previous work [17, 15], in which:
  1. 1.
    Maximising simplicity of policies was proven unnecessary and insufficient to maximise sample efficiency [17, prop. 3].
  2. 2.
    Maximising policy “weakness” was proven necessary and sufficient to maximise sample efficiency [17, prop. 1,2] and identify cause [15]. In experiments, weak policies outperformed simple by $110-500\%$.
Our purpose here is to extend this work, and to establish:
  1. 1.
    Is complexity just an artefact of abstraction? Does it measure anything meaningful in the absence of abstraction?
  2. 2.
    Why does there tend to be a correlation between sample efficiency and simplicity?
1.0.3 Results:
Our results are only meaningful if one accepts that this formalism is reflective of reality, and so to begin we provide an argument to the effect that it is (lest we be accused of straw-manning complexity). We also heavily modify and extend and refine the formalism to this end, so the formalism itself may be considered a result (we change nothing that could affect the outcome of previously published proofs). Second, we show that the complexity of all behaviours is equal in the absence of an abstraction layer (a general formalisation of any interpreter). In other words, complexity is a subjective “illusion”. We further show that if the vocabulary is finite then weakness can confound simplicity and sample efficiency. Third, we argue that abstraction is goal directed. If the vocabulary is finite and tasks are uniformly distributed, then efficiency demands that weak statements take simple forms, causing the confounding.
2 The Formalism
Definition 1 (environment)
  • We assume an infinite set $\Phi$ whose elements we call states.
  • A declarative program is $f\subseteq\Phi$, and we write $P$ for the set of all declarative programs (the powerset of $\Phi$).
  • By a truth or fact about a state $\phi$, we mean $f\in P$ such that $\phi\in f$.
  • By an aspect of a state $\phi$ we mean a set $l$ of facts about $\phi$ s.t. $\phi\in\bigcap l$. By an aspect of the environment we mean an aspect $l$ of any state, s.t. $\bigcap l\neq\emptyset$. We say an aspect of the environment is realised by state $\phi$ if it is an aspect of $\phi$.
Definition 2 (abstraction layer)
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite unless explicitly stated otherwise. If $\mathfrak{v}=P$, then we say that there is no abstraction.
  • $L_{\mathfrak{v}}=\{l\subseteq\mathfrak{v}:\bigcap l\neq\emptyset\}$ is a set of aspects in $\mathfrak{v}$. We call $L_{\mathfrak{v}}$ a formal language, and $l\in L_{\mathfrak{v}}$ a statement.
  • We say a statement is true iff it is an aspect of the environment realised by the present state.
  • A completion of a statement $a$ is a statement $b$ which is a superset of $a$. If $b$ is true, then $a$ is true.
  • The extension of an statement $a\in{L_{\mathfrak{v}}}$ is $E_{a}=\{b\in{L_{\mathfrak{v}}}:a\subseteq b\}$. $E_{a}$ is the set of all completions of $a$.
  • The extension of a set of statements $A\subseteq{L_{\mathfrak{v}}}$ is $E_{A}=\bigcup\limits_{a\in A}E_{a}$.
  • We say $a$ and $b$ are equivalent iff $E_{a}=E_{b}$.
(notation) $E$ with a subscript is the extension of the subscript111e.g. $E_{l}$ is the extension of $l$..
(intuitive summary) $L_{\mathfrak{v}}$ is everything which can be realised in this abstraction layer. The extension $E_{a}$ of a statement $a$ is the set of all statements whose existence implies $a$, and so it is like a truth table.
Definition 3 ($\mathfrak{v}$-task)
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
  • ${I}_{\alpha}\subset L_{\mathfrak{v}}$ is a set whose elements we call inputs of $\alpha$.
  • ${O_{\alpha}}\subset E_{I_{\alpha}}$ is a set whose elements we call correct outputs of $\alpha$.
${I_{\alpha}}$ has the extension $E_{I_{\alpha}}$ we call outputs, and ${O_{\alpha}}$ are outputs deemed correct. $\Gamma_{\mathfrak{v}}$ is the set of all tasks given $\mathfrak{v}$.
[…]
(notation) If $\omega\in\Gamma_{\mathfrak{v}}$, then we will use subscript $\omega$ to signify parts of $\omega$, meaning one should assume $\omega=\langle{I}_{\omega},{O}_{\omega}\rangle$ even if that isn’t written.
(intuitive summary) To reiterate and summarise the above:
  • An input is a possibly incomplete description of a world.
  • An output is a completion of an input [def. 2].
  • A correct output is a correct completion of an input.
2.0.1 Learning and inference definitions:
Machine learning is typically divided into learning and inference. To learn a task, an organism learns a policy (strictly speaking, an organism is a policy). By constraining how we complete inputs, a policy allows us to do abductive inference (much as a constraint in a SAT solver would). Inference requires a policy and learning a policy requires a proxy, the definitions of which follow.
Definition 4 (inference)
  • A $\mathfrak{v}$-task policy is a statement $\pi\in L_{\mathfrak{v}}$. It constrains how we complete inputs.
  • $\pi$ is a correct policy iff the correct outputs $O_{\alpha}$ of $\alpha$ are exactly the completions $\pi^{\prime}$ of $\pi$ such that $\pi^{\prime}$ is also a completion of an input.
  • The set of all correct policies for a task $\alpha$ is denoted $\Pi_{\alpha}$. [^2] To repeat the above definition in set builder notation: ${\Pi_{\alpha}}=\{\pi\in L_{\mathfrak{v}}:{E}_{I_{\alpha}}\cap E_{\pi}={O_{\alpha}}\}$
Assume $\mathfrak{v}$-task $\omega$ and a policy $\pi\in L_{\mathfrak{v}}$. Inference proceeds as follows:
  1. 1.
    we are presented with an input ${i}\in{I}_{\omega}$, and
  2. 2.
    we must select an output $e\in E_{i}\cap E_{\pi}$.
  3. 3.
    If $e\in{O}_{\omega}$, then $e$ is correct and the task “complete”. $\pi\in{\Pi}_{\omega}$ implies $e\in{O}_{\omega}$, but $e\in{O}_{\omega}$ doesn’t imply $\pi\in{\Pi}_{\omega}$.
(intuitive summary) To reiterate and summarise the above:
  • A policy constrains how we complete inputs.
  • A correct policy is one that constrains us to correct outputs.
In functionalist terms, a policy is a “causal intermediary”.
Definition 5 (learning)
  • A proxy $<$ is a binary relation on statements, and the set of all proxies is $Q$.
  • $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$.
  • $<_{d}$ is the description length or simplicity proxy. We have $l_{1}<_{d}l_{2}$ iff $\lvert l_{1}\rvert>\lvert l_{2}\rvert$.
By the weakness of a statement, we mean the cardinality of its extension. By the weakness of an extension we mean its cardinality. Likewise, when we speak of simplicity with regards to a statement, we mean its cardinality. The complexity of an extension is the simplest statement of which it is an extension.
(generalisation) A statement $l$ generalises to a $\mathfrak{v}$-task $\alpha$ iff $l\in\Pi_{\alpha}$. We speak of learning $\omega$ from $\alpha$ iff, given a proxy $<$, $\pi\in{\Pi}_{\alpha}$ maximises $<$ relative to all other policies in ${\Pi}_{\alpha}$, and $\pi\in{\Pi}_{\omega}$.
(probability of generalisation) We assume a uniform distribution over $\Gamma_{\mathfrak{v}}$. If $l_{1}$ and $l_{2}$ are policies, we say it is less probable that $l_{1}$ generalizes than that $l_{2}$ generalizes, written $l_{1}<_{g}l_{2}$, iff, when a task $\alpha$ is chosen at random from $\Gamma_{\mathfrak{v}}$ (using a uniform distribution) then the probability that $l_{1}$ generalizes to alpha is less than the probability that $l_{2}$ generalizes to $\alpha$.
(sample efficiency) Suppose $\mathfrak{app}$ is the set of all pairs of policies. Assume a proxy $<$ returns $1$ iff true, else 00. Proxy $<_{a}$ is more sample efficient than $<_{b}$ iff
$\left(\sum_{(l_{1},l_{2})\in\mathfrak{app}}\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{a}l_{2})\rvert-\lvert(l_{1}<_{g}l_{2})-(l_{1}<_{b}l_{2})\rvert\right)<0$
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [17]).
(intuitive summary) Learning is an activity undertaken by some manner of intelligent agent, and a task has been “learned” by an agent that knows a correct policy. Humans typically learn from “examples”. An example of a task is a correct output and input. A collection of examples is a child task, so “learning” is an attempt to generalise from a child one of its parents. The lower level the child from which an agent generalises to parent, the “faster” it learns, the more sample efficient the proxy. The most sample efficient proxy is $<_{w}$.
3 Arguments and Results
The distinction between software and hardware, though useful, is unsuitable for reasoning about cause [15]. A software agent in an interactive setting performs subjectively, depending on hardware through which it is interpreted [14]. If we are to understand complexity, we must understand what the concept entails in the absence of such “abstraction layers”. We must ascertain what is objective rather than subjective, and so we must begin at the level of the environment rather than the agent. To this end, previous work [18, 17, 15, 19] proposed a “de facto” pancomputational [20] model of all possible environments and aspects thereof, including all abstractions and goal directed behaviours, which we refine and justify for here. Our formalism is not computational in the sense of dealing with symbols and quantities, or any other high level abstraction interpreted by a human mind. The assumptions we make are extremely weak.
Axiom 1: There are things which exist, and things which do not. We call the things which exist the environment.
Axiom 2: The environment has at least one state. To say it has more than one state is to claim there exists at least one “dimension”. A dimension is a set of points, for example time, we do not make the additional claim that there is any order to that set. Each state is the environment at a different point. Each state is “extended” along a dimension in the same sense as material or mental substance, meaning no two states can occupy the same points in all dimensions (in other words, states are like the environment perceived from different positions in time or some other dimensions by an omniscient observer).
These are all the assumptions we need for our version of pancomputationalism. We don’t even need to speculate about any internal structure states might have. States differ along one or more dimensions. Any “fact” about a state’s internal structure can be defined by its relation to other states. The existence of sets is implied by the existence of states (because there is more than one “thing” that exists), and a fact is just the set of states in which it holds (the truth conditions of a “declarative program”).
Universality Claim: Axioms 1 and 2 hold for every conceivable environment.
As truth is defined in existential terms, there is nothing which is not a fact of the environment, and no environment which we cannot describe as a set of facts. In other words, this is a minimalist formalism of everything. These are not statements evaluated by a mind, and truth is not semantic in that sense. An environment could be like our own, or not. It could be deterministic, in which case states follow a sequence, or non-deterministic, in which case they don’t [^3] In any case, the difference between deterministic and non-deterministic seems meaningless when you consider that DFAs and NFAs are equivalent [21].. It could be fantastical with magic and true names. It could even be a world constructed through the subjective experience of its inhabitants. All that matters is that an environment has states differentiated along one or more dimension, and from there we obtain the set of all facts. This gives us some notion of internal structure. An aspect of a state is just a set of facts about that state. With aspects in hand, we can define abstraction. An aspect is akin to a logical statement. It has a truth value given a state. We also define its extension (def. 2), which is all other aspects of which it is a part. This implies a “lattice” of aspects, and the existence of abstraction. An abstraction layer is just a set of declarative programs, called a vocabulary, which acts as a formal language whose rules are determined by states. An abstraction layer implies a set of “$\mathfrak{v}$-tasks” (def. 3), each of which is behaviour that defines a system. The formalisation of policies as causal intermediaries between inputs and outputs [22] then develops this into a causal depiction of goal directed behaviour (assuming a policy is implied by the inputs and outputs). Again, we must emphasise this is a first principles approach. We do not assume symbols and Turing machines. Inputs, outputs and policies are all just sets of declarative programs. Whether something is goal directed is determined by the relations between states. Finally, we must now define complexity. The claims we make in the rest of this paper pertain to this notion of complexity.
Complexity of an Behaviour: If two statements have the same extension, then the simpler is the smaller set (def. 5). The complexity of an extension is the simplest statement of which it is an extension. This is analogous to other measures of complexity [2, 3], but facilitates comparison of abstraction layers.
3.1 Implications for Complexity
Proposition 1 (subjectivity)
If there is no abstraction, complexity can always be minimized without improving sample efficiency, regardless of the task [^4] In other words, complexity has no influence on sample efficiency..
Proof
In accord with definition 2, the absence of abstraction means the vocabulary is the set of all declarative programs, meaning $\mathfrak{v}=P$. It follows that for every $l\in L_{\mathfrak{v}}$ there exists $f\in\mathfrak{v}$ such the $\bigcap l=\{f\}$. Statements $l$ and $\{f\}$ are equivalent iff $E_{l}=E_{\{f\}}$, which is exactly the case here because $\bigcap l=\{f\}$. [17, prop. 1,2] shows maximising weakness is necessary and sufficient to maximise the probability of generalisation, which means weakness maximises sample efficiency (is the optimal proxy). This means sample efficiency is determined by the cardinality of extension. For every correct policy $l$ of every task in $\Gamma_{\mathfrak{v}}$ there exists $f\in\mathfrak{v}$ s.t. $E_{l}=E_{\{f\}}$. Policy complexity can be minimised regardless weakness, because minimum description length representation of every extension is $1$.
In this sense, complexity is an illusion created by abstraction. It is meaningless, because all behaviours (extensions) are implied by statements of the same complexity. However, that doesn’t change the fact that simpler forms have been observed to generalise better. This raises the question; what additional assumptions can we make that would explain the correlation?
3.1.1 Time, space and causal confounding:
We now make the additional assumption that vocabularies are finite. Every aspect of our environment appears to be spatially and temporally extended, so for the sake of understanding complexity we make this assumption for all environments [^5] This something is spatially and temporally extended if it occupies a location in space and time, and nothing else can occupy that same location.. We hold that this justifies the assumption of a finite vocabulary, because in our spatially and temporally extended environment the amount of information which can influence a system in a finite area of time and space is finite [23].
Proposition 2 (confounding)
If the vocabulary is finite, then policy weakness can confound [^6] Confounding [24] is when $A=\mbox{``{\tt patient older}''}$ ” causes $B=\mbox{``{\tt died from covid}''}$ ” and $C=\mbox{``{\tt given vaccine}''}$ ”, but $C$ and $B$ have no influence over each other and so it might appear like $C$ causes $B$, because $p(B\mid C)>p(B\mid\lnot C)$, but the underlying cause becomes apparent when we see $p(B,C\mid A)<p(B,\lnot C\mid A)$. sample efficiency with policy simplicity.
Proof
We already have that policy weakness causes sample efficiency, in that it is necessary and sufficient to maximise it in order to maximise sample efficiency. Continuing from proof 1, in a finite vocabulary, there may not exist $f\in\mathfrak{v}$ s.t. $E_{l}=E_{\{f\}}$, which means the complexity of all extensions will not be the same. If we choose any vocabulary in which weaker aspects take simpler forms, then simplicity will be correlated with weakness and so will also be correlated with sample efficiency. This means we would choose $\mathfrak{v}$ s.t. for all $a,b\in L_{\mathfrak{v}}$, the simpler statement has the larger extension, meaning $a<_{w}b\leftrightarrow a<_{d}b$.
3.1.2 Why confounding tends to occur:
We now briefly argue that abstraction is goal directed, meaning the tasks an abstraction layer tends to represent are those it is best suited to represent, and that this implies weak constraints take simple forms. Goal directedness is formalised as a $\mathfrak{v}$-task, so one way we might formalise goal directed abstraction is as a mapping from $\Gamma_{\mathfrak{v}}$ to abstraction layers (or more specifically vocabularies, which are subsets of $P$). There is an equivalent declarative program for every aspect, meaning no matter the abstraction layer for every statement $l$ there exists $f\in P$ s.t. if $f\in\mathfrak{v}$ then $E_{l}=E_{\{f\}}$, the outputs of a $\mathfrak{v}$-task are equivalent to a new vocabulary $\mathfrak{v^{\prime}}$, which implies an abstraction layer in which $\mathfrak{v^{\prime}}$-tasks can be expressed. There are several reasons to believe the formation of abstraction layers is goal directed, depending upon the context in which we consider complexity. In the case of a computer, a human has specifically designed each abstraction layer to express that which is needed for a purpose. What separates x86 from a higher level abstraction layer like Numpy is that the former has a more general intended purpose, expressing “weaker” constraints. We tend to construct abstraction layers to be as versatile as possible whilst satisfying a particular need. More generally natural selection favours adaptation, which means generalisation, which is maximised by preferring weaker policies [17]. Multiscale competency architectures observed in living organisms [25, 26] amount to self organising abstraction layers. Because natural selection favours adaptable organisms, these abstraction layers must adapt to represent the weakest policies which constitute fit behaviour. More generally, we speculate that phase transitions motivate the emergence of self preserving goal directed behaviour, by destroying some physical structures and preserving others. Such goal directed abstraction must minimise the size of vocabularies at higher levels, whilst also maximising the weakness of the policies they can express (two opposing pressures). This is because a larger a vocabulary exponentially increases the space of outputs and policies [27], which may conflict with with finite time and space constraints. A larger vocabulary would make inference and learning less tractable (more “complex” in the sense of being a more difficult search problem that takes up more time). To maximise the weakness of policies in higher levels of abstraction, while minimising the size of the vocabulary in which they’re expressed, weaker policies must take simpler forms.